OpenStudy (anonymous):

-x^2 + 3 = (2x - 1)^2 What is this quadratic equation written in standard form?

OpenStudy (jdoe0001):

expand the binomial and cancel out like-terms and add

OpenStudy (anonymous):

Using what @jdoe0001 said.... -x^2 + 3 = (2x - 1)^2 -x^2 + 3 = 4 x^2-4 x+1 -5 x^2+4 x+2 = 0